1. Microwave ovens emit microwave energy with a wavelength of 12.5 cm. What is the energy of exactly one photon of this microwave radiation?
2. A certain shade of blue has a frequency of 7.18x10^14 Hz. What is the energy of exactly one photon of this light?
Answer #1

Ans1= we know that E=hc/lemda where

h= Planck’s Constant = 6.626×10^(-34) J∙s

  c= speed of light = 3.00×10^8 m

λ=12.5cm= 0.125 m

E=( (6.626×10^(-34) J∙s)(3.00×10^8 m/s) ) / (0.125 m )=1.59×10^(-24) J